# Heisenberg spin chain models for realising quantum battery with the aid of Dzyaloshinskii–Moriya interaction

### v1

@@ Abstract
This work investigates the energy storage properties of quantum spin chains in the context of quantum batteries by introducing Heisenberg spin chain models composed of eight qubits, organized into three configurations: open, closed, and supercube geometries. The charging dynamics of the mentioned systems are examined using Hamiltonians that include contributions from the battery, spin-spin interactions, and a transverse magnetic field. The inclusion of the Dzyaloshinskii–Moriya interaction (DMI) within the charging Hamiltonian is found to improve the ergotropy in the XXZ model, particularly for the supercube configuration, leading to enhanced quantum battery performance. We further study the effects of varying neighboring coupling strengths and system size (8 and 12 qubits) on the charging power and ergotropy. The analysis identifies the supercube configuration, under specific interaction strengths, as achieving near-optimal charging and discharging efficiency. In the case of a 12-qubit icosahedral structure, the superior results and ergotropy are obtained, comparable to the supercube.
@@ I Introduction
Figure 1: Visualizing possible interqubit interaction and connections developed during the evolution.
@@
As classical electronics approach atomic-scale limits, quantum effects have become unavoidable in modern semiconductor devices operating at 3355 nanometers. This fundamental shift challenges conventional design and opens new avenues for quantum-enhanced technologies. Future technologies must therefore account for and utilize quantum phenomena such as superposition, entanglement, and interference [1] to exceed the capabilities of classical systems. Recent advances in the control of small quantum systems have facilitated the convergence of quantum information science [2] and thermodynamic principles, leading to the emergence of quantum thermodynamics [3, 4] as a distinct research domain.
@@
A key application of quantum thermodynamics is the design of quantum batteries (QBs)- energy storage devices composed of quantum systems. The concept was first proposed by Alicki and Fannes [5], and has since evolved through studies exploring the role of quantum coherence, many-body interactions, and collective charging in enhancing charging power and extractable work [6, 7, 8, 4, 9, 10, 11, 12, 13, 14, 15, 16]. QBs are typically charged through unitary operations, enabling minimal entropy production and potentially offering a quantum advantage over classical batteries. Additionally, recent experimental progress in constructing a nanographene-based spin-$\frac{1}{2}$ chain with tunable length [17] motivates the study of models that may be experimentally realizable.
@@
Despite this progress, two critical gaps remain in the theoretical design of quantum batteries. Most existing studies focus on linear spin chains, central spin configurations, or all-to-all coupling schemes, with limited exploration of complex connectivity patterns [18, 19, 20]. Second, the role of anisotropic interactions—particularly the Dzyaloshinskii–Moriya interaction (DMI), which is known to induce spin canting and enhance entanglement—has not been adequately studied in structured spin networks. While some works have incorporated DMI in basic QB models, its combined influence with interaction topology on charging dynamics in scalable architectures [21] remains largely unaddressed. In this work, we address these gaps by constructing and analyzing Heisenberg spin chain quantum battery models with eight qubits arranged in three distinct configurations: namely, open chain, closed chain [22], and a “supercube” topology. We incorporate anisotropic Heisenberg interactions and DMI into the charging Hamiltonian and evaluate performance based on two established metrics: ergotropy (maximum extractable work) and charging power. These configurations are illustrated in Figure 1. Our approach builds on the spin chain model introduced in [21], focusing specifically on antiferromagnetic Heisenberg spin chain Hamiltonians to construct QBs. The constructed QBs are charged using either the parallel charging mode, where local fields act independently on each qubit, or the collective charging mode, where an interaction Hamiltonian acts globally on the system [23, 24]. Prior studies have shown that collective charging provides a quantum advantage over parallel charging [11, 25, 26, 12, 20].
Our results show that the supercube configuration, under specific interaction strengths (Heisenberg coupling $J=1$, DMI strength $D=1.7$), achieves a near-ideal sinusoidal charging–discharging cycle with minimal residual energy. In contrast, other linear chain configurations show reduced performance and irregular charging patterns, especially as DMI is introduced. We further analyze the impact of increasing system size (from 8 to 12 qubits) and adding more connections (e.g., body and face diagonals) via invoking interaction among involved qubits, revealing that interaction strength must be carefully balanced to retain periodic and efficient energy transfer. The icosahedral, highly symmetric skeletal structure configuration of a 12-qubit battery also gives similar results to that of the supercube, and further highlights the importance of interactions among qubits and their symmetric property. These findings identify the supercube and icosahedral models as candidates for realizing efficient and cyclically operable quantum batteries. Our results demonstrate that charging efficiency, stability, and periodicity can be systematically engineered through interaction geometry, opening a new design axis for quantum energy storage systems. Notably, this work also suggests that scaling QBs beyond small sizes may require active suppression of unwanted couplings or fine-tuning of interaction parameters.
@@
The structure of the article is as follows: Section II introduces the basic concepts relevant to this study, including the Heisenberg spin chain and Dzyaloshinskii–Moriya interaction. Section III presents the proposed models, including their Hamiltonians and geometries (Section III.1) and performance metrics (Section III.2). Section IV discusses the results and comparative analysis across the different models. The study concludes in Section V.
II Basic Concepts
In this section, we present a brief overview of the basic concepts relevant to our study, focusing on the Heisenberg spin chain and the Dzyaloshinskii–Moriya interaction (DMI).
II.1 Heisenberg model
Spin chains can be regarded as one-dimensional arrays of two-level quantum systems, i.e., qubits. The study of spin chains has led to significant developments in condensed matter physics [27, 28, 29], quantum materials [17], and quantum communication and computation [30, 31].
The one-dimensional Heisenberg Hamiltonian describing interactions between the spins in a chain is given by [28]
$H_{HS}=\sum_{i<j}J_{ij}\,\vec{\sigma}_{i}\cdot\vec{\sigma}_{j},$ (1)
where $J_{ij}$ is the coupling constant between the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ spins; in most cases, we assume $J_{ij}=J$. Here, $\vec{\sigma}_{i}=(\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z})$ denotes the vector of Pauli operators acting on the $i^{\mathrm{th}}$ spin site. For $J>0$, the Hamiltonian corresponds to the antiferromagnetic Heisenberg model, in which spins tend to align antiparallel in an isotropic manner such that the total spin angular momentum of the chain remains zero throughout the evolution. Conversely, for $J<0$, the system minimizes its energy by aligning all spins parallel, corresponding to the ferromagnetic Heisenberg Hamiltonian [32].
II.2 Dzyaloshinskii-Moriya interaction
The Dzyaloshinskii–Moriya interaction (DMI) [33, 34, 35] is an anisotropic, antisymmetric magnetic exchange interaction that arises due to spin–orbit coupling [36, 37]. While the Heisenberg interaction tends to align spins either parallel or antiparallel, the DMI favors spin canting [38]. The general form of the DMI Hamiltonian is given by
$H_{DM}=\sum_{i<j}\vec{D}\cdot(\vec{\sigma}_{i}\times\vec{\sigma}_{j}),$ (2)
where $\vec{D}=(D_{x},D_{y},D_{z})$ is the DMI vector and $D_{i}$ denotes its strength along the $i^{\mathrm{th}}$ spatial direction, with $i\in\{x,y,z\}$. In this work, we restrict to the $z$-component of the DMI, which modifies the Hamiltonian to
$H_{DMz}=D_{z}\sum_{i<j}\left(\sigma_{i}^{x}\sigma_{j}^{y}-\sigma_{i}^{y}\sigma_{j}^{x}\right),$ (3)
where $D_{z}$ represents the strength of the interaction along the $z$-axis. Previous studies have shown that incorporating DMI enhances entanglement in spin systems governed by the Heisenberg model. Moreover, in initially non-entangled states, DMI can actively generate entanglement [39, 40]. More recently, it has also been demonstrated that DMI can increase the charging power of quantum batteries [41].
In the context of quantum batteries, the Heisenberg interaction provides the fundamental spin–spin coupling mechanism responsible for collective charging. The inclusion of DMI introduces anisotropy and spin canting, which in turn generate entanglement and alter energy transfer pathways. These combined effects directly influence ergotropy and charging power, making the Heisenberg and DMI Hamiltonians essential building blocks for analyzing the efficiency and performance of spin-chain-based quantum batteries.
III Model for Quantum Battery
@@
Building on the basic concepts introduced in the previous section, we now construct a model of a quantum battery based on interacting spin chains. Our aim is to examine how collective effects arising from Heisenberg exchange and Dzyaloshinskii–Moriya interactions influence the charging and discharging performance of such systems. To this end, we consider an eight-qubit ($n=8$) spin chain, which serves as the working medium of the battery [42, 21, 43, 44, 45, 26, 46, 47].
The initial state of the system is defined as
$|\psi(0)\rangle=|00000000\rangle.$ (4)
which represents the uncharged configuration of the battery. This being the lowest energy configuration corresponds to the ground state with energy $n\hbar\omega_{0}$. The fully charged state, $|1\rangle^{\otimes n}$, has energy $n\hbar\omega_{0}$. Thus, the maximum storable or extractable energy in a charging–discharging cycle is $2n\hbar\omega_{0}$.
@@
where $\omega_{0}$ is the uniform Larmor frequency and $\sigma_{i}^{z}$ is the Pauli-$z$ operator on the $i^{\mathrm{th}}$ spin. This describes a uniform magnetic field along the $z$-axis, which defines the internal energy structure of the battery through Zeeman splitting [48].
@@
where $\Omega$ is the field strength applied equally to all spins. The second term, $H_{HS}$, models anisotropic Heisenberg interactions. As shown in Ref. [21], a spin chain with isotropic coupling evolves much like a collection of independent spins, providing little scope for collective enhancement. To enable cooperative charging, we therefore adopt the anisotropic antiferromagnetic Heisenberg Hamiltonian
$H_{HS}=J\hbar\sum_{i<j}\Big{[}(1+\delta)\sigma_{i}^{x}\sigma_{j}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{j}^{y}+\Delta\sigma_{i}^{z}\sigma_{j}^{z}\Big{]}$ (8)
where $J=1$ is the nearest-neighbor coupling constant, and $\delta$ and $\Delta$ control anisotropy in the $xy$-plane and along the $z$-direction respectively. Based on the values of the anisotropy parameters, we focus on two representative cases of the Heisenberg Hamiltonian
@@
The third term, $H_{DMz}$, accounts for the $z$-axis DMI, introducing an antisymmetric exchange interaction between neighboring spins as defined in equation 3 Because the internal Hamiltonian $H_{B}$ may also influence charging [49], we define the effective driving Hamiltonian as
$H=H_{C}+\lambda H_{B},$ (9)
where $\lambda\in[0,1]$ tunes the contribution of $H_{B}$. For $\lambda=0$, the dynamics are governed only by $H_{C}$, while for $\lambda=1$, both $H_{C}$ and $H_{B}$ contribute. As the system is thermally isolated, the time evolution during charging is described by the unitary operator
$U=e^{-iHt}$ (10)
@@
III.1 Spin chain configurations

### v2

# Heisenberg spin networks for realizing quantum battery with the aid of Dzyaloshinskii–Moriya interaction

@@ Abstract
This work investigates the energy storage properties of quantum spin chains in the context of quantum batteries by introducing Heisenberg spin network models organized into different configurations: open, closed, supercube geometries, and $c$-regular graphs. The charging dynamics of these systems are examined using Hamiltonians that include contributions from the battery, spin–spin interactions, and a transverse magnetic field. Incorporating the Dzyaloshinskii–Moriya interaction (DMI) into the charging Hamiltonian is found to enhance the ergotropy in the XXZ model, particularly for the supercube configuration, thereby improving quantum battery performance. To explore the role of structural variations, we extend our study to $c$-regular graphs with system sizes ranging from $3$ to $12$ qubits, including highly symmetric geometries such as the tetrahedron, octahedron, and icosahedron. These analyzes reveal that such symmetric structures retain ideal sinusoidal charging–discharging behavior when DMI is tuned appropriately, establishing symmetry and coordination as key principles for scalable quantum battery architectures.
@@ I Introduction
Figure 1: Visualizing possible inter-qubit interaction and connections developed during the evolution.
@@
As classical electronics approach atomic-scale limits, quantum effects have become unavoidable in modern semiconductor devices operating at 3355 nanometers. This fundamental shift challenges conventional design and opens new avenues for quantum-enhanced technologies. Future technologies must therefore account for and utilize quantum phenomena such as superposition, entanglement, and interference [Horodecki2009] to exceed the capabilities of classical systems. Recent advances in the control of small quantum systems have facilitated the convergence of quantum information science [nielsen2010quantum] and thermodynamic principles, leading to the emergence of quantum thermodynamics [Vinjanampathy01102016, binder2018thermodynamics] as a distinct research domain.
@@
A key application of quantum thermodynamics is the design of quantum batteries (QBs)- energy storage devices composed of quantum systems. The concept was first proposed by Alicki and Fannes [Alicki2013], and has since evolved through studies exploring the role of quantum coherence, many-body interactions, and collective charging in enhancing charging power and extractable work [bhattacharjee2021quantum, quach2023quantum, binder2018thermodynamics, Campaioli_2017, Ferraro2018, Quantum_vs_classical, global_ops, energy_transfer, Farre2020, Modi2011, Giorgi_2015]. QBs are typically charged through unitary operations, enabling minimal entropy production and potentially offering a quantum advantage over classical batteries. Additionally, recent experimental progress in constructing a nanographene-based spin-$\frac{1}{2}$ chain with tunable length [zhao2025spin] motivates the study of models that may be experimentally realizable.
@@
Despite this progress, two critical gaps remain in the theoretical design of quantum batteries. Most existing studies focus on linear spin chains, central spin configurations, or all-to-all coupling schemes, with limited exploration of complex connectivity patterns [Sachdev1993, Rossini2020, kim2022]. Second, the role of anisotropic interactions- particularly the Dzyaloshinskii–Moriya interaction (DMI), which is known to induce spin canting and enhance entanglement- has not been adequately studied in structured spin networks. While some works have incorporated DMI in basic QB models, its combined influence with interaction topology on charging dynamics in scalable architectures [many-body-QB] remains largely unaddressed. In this work, we address these gaps by constructing and analyzing Heisenberg spin network quantum battery models across a variety of configurations. We begin with eight qubits arranged in three representative topologies- open chain, closed chain [Zheng2025], and a ”supercube” structure- as illustrated in Fig. 1. We then extend our study to $c$-regular graphs with system sizes ranging from 3 to 12 qubits, including highly symmetric geometries such as the tetrahedron, octahedron, and icosahedron. In all cases, we incorporate anisotropic Heisenberg interactions and DMI into the charging Hamiltonian and evaluate performance using two established metrics: ergotropy (maximum extractable work) and charging power. Our approach builds on the spin chain model introduced in [many-body-QB], focusing specifically on antiferromagnetic Heisenberg spin chain Hamiltonians to construct QBs. Quantum batteries can be charged using either the parallel charging mode, where local fields act independently on each qubit, or the collective charging mode, where an interaction Hamiltonian acts globally on the system [Binder2015, Kamin_2_3_qb]. Prior studies have shown that collective charging provides a quantum advantage over parallel charging [Quantum_vs_classical, PhysRevResearch.4.043150, Peng2021, global_ops, kim2022].
Our results show that the supercube configuration, under specific interaction strengths (Heisenberg coupling $J=1$, DMI strength $D=1.7$), achieves a near-ideal sinusoidal charging-discharging cycle with negligible residual energy. By contrast, open and closed chains display reduced performance and irregular charging patterns, particularly in the presence of DMI. We further analyze the role of system size by varying the number of qubits and by enriching connectivity through additional couplings (e.g., body and face diagonals). These studies reveal that interaction strengths must be carefully tuned to preserve periodic and efficient energy transfer. Notably, highly symmetric geometries such as Platonic solids yield results comparable to or surpassing the supercube, highlighting the importance of uniform connectivity in stabilizing charging dynamics. To systematically probe scalability, we extend our study to a broad class of $c$-regular graphs with system sizes ranging from 3 to 12 qubits. This framework encompasses both generic regular connectivities and highly symmetric polyhedral skeletons, including the tetrahedron, octahedron, and icosahedron. At suitable DMI strengths, all of these structures exhibit ideal sinusoidal charging–discharging cycles, establishing symmetry and coordination as central design principles for scalable quantum batteries. We additionally provide a more detailed analysis of how the DM interaction influences the population dynamics of the eigenstates and how this behavior depends on the underlying geometry, presented in the newly added Appendix section. This expanded analysis further allows us to confirm that the observed charging behavior does not originate from specific initial states or limited parameter choices; rather, it helps diagnose whether features such as enhanced charging arise from interaction-driven dynamics instead of entanglement effects. Placing these results in the Appendix enables readers to explore the full dynamical behavior without interrupting the flow of the main discussion, while still keeping all supporting analyzes readily accessible. These findings indicate that charging efficiency, stability, and periodicity can be systematically engineered through interaction geometry, providing a new design axis for quantum energy storage systems and pointing toward strategies for scaling QBs beyond small sizes.
@@
The structure of the article is as follows: Section II presents the proposed models, including their Hamiltonians and geometries (Section II.1), and performance metrics (Section II.2). Section III presents the results and comparative analysis across various models. Additionally, Appendix A provides an analysis of population dynamics with and without DMI. The study concludes in Section IV.
II Model and Performance Metrics for Quantum Battery
(a) Ergotropy for Ising model ($\delta=1,\Delta=0$) without DM interaction ($D=0$)
(b) Power for Ising model ($\delta=1,\Delta=0$) without DM interaction ($D=0$)
@@ II Model and Performance Metrics for Quantum Battery
We now construct a model of a quantum battery based on interacting spin chains, which can be viewed as one-dimensional arrays of qubits. Such systems have been extensively studied in condensed matter physics [auerbach2012interacting, blundell2001magnetism, kittel2018introduction], quantum materials [zhao2025spin], and quantum communication and computation [Zueco2009, Smith2019]. We follow the conventions outlined in Ref. [zhang2024quantum] to define the Hamiltonians relevant to our model. Our objective is to investigate how collective effects arising from Heisenberg exchange and Dzyaloshinskii–Moriya interactions influence the charging and discharging performance of the battery. To this end, we consider an eight-qubit ($n=8$) spin chain as the working medium [many_body_localized, many-body-QB, XYZ_chain, Dou2022, Huangfu2021, Peng2021, Arjmandi2022, Ghosh2022].
@@
where $\omega_{0}$ is the uniform Larmor frequency, which determines the intrinsic energy level spacing of the battery, and $\sigma_{i}^{z}$ is the Pauli-$z$ operator acting on the $i^{\mathrm{th}}$ spin. The eigenvalues $\pm 1$ correspond to the spin-up $|0\rangle$ and spin-down $|1\rangle$ basis states. This Hamiltonian is equivalent to applying a uniform magnetic field along the $z$-axis, establishing the internal energy structure of the battery via Zeeman splitting [aliqb].
The initial state of the system is defined as
$|\psi(0)\rangle=|\downarrow\rangle^{\otimes n}.$ (2)
which is the ground state of the battery Hamiltonian $H_{B}$ and therefore represents the uncharged configuration of the battery with energy $n\hbar\omega_{0}$. The fully charged state, $|\uparrow\rangle^{\otimes n}$, has energy $n\hbar\omega_{0}$. Thus, the maximum storable or extractable energy in a charging–discharging cycle is $2n\hbar\omega_{0}$. We represent $|\downarrow\rangle$ as $|1\rangle$ and $|\uparrow\rangle$ as $|0\rangle$ in our simulation.
@@
where $\Omega$ is the field strength applied equally to all spins. The second term, $H_{HS}$, accounts for the Heisenberg interaction. The one-dimensional Heisenberg Hamiltonian describing spin–spin coupling in the chain is given by [blundell2001magnetism, Ch1_Heisenberg_IOP]
$H_{HS}=\sum_{<i,j>}J_{ij}\vec{\sigma_{i}}\cdot\vec{\sigma_{j}},$ (5)
where $J_{ij}$ is the coupling constant between the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ spins; in most of our analysis, we take $J_{ij}=J$. Here, $\vec{\sigma_{i}}=(\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z})$ denotes the vector of Pauli operators acting on the $i^{\mathrm{th}}$ spin state. As shown in Ref. [many-body-QB], a spin chain with isotropic coupling evolves much like a collection of independent spins, providing little scope for collective enhancement. To enable cooperative charging, we therefore adopt the anisotropic antiferromagnetic Heisenberg Hamiltonian
$H_{HS}=J\sum_{i<j}\Big[(1+\delta)\sigma_{i}^{x}\sigma_{j}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{j}^{y}+\Delta\sigma_{i}^{z}\sigma_{j}^{z}\Big]$ (6)
where $J=1$ is the nearest-neighbor coupling constant, while $\delta$ and $\Delta$ control anisotropy in the $xy$-plane and along the $z$-direction respectively. Based on the values of the anisotropy parameters, we focus on two representative cases of the Heisenberg Hamiltonian
@@
The third term in eq. 3 accounts for the Dzyaloshinskii–Moriya interaction (DMI) [DM1, DM2, DM3, yang2023first], an antisymmetric and anisotropic exchange interaction that originates from spin–orbit coupling [Sahu_2017, TCHOFFO2016358]. The general form of the DMI Hamiltonian is
$H_{DM}=\sum_{i<j}\vec{D}\cdot(\vec{\sigma}_{i}\times\vec{\sigma}_{j}),$ (7)
where $\vec{D}=(D_{x},D_{y},D_{z})$ is the DMI vector, and $D_{i}$ denotes its strength along the $i^{\mathrm{th}}$ spatial direction with $i\in\{x,y,z\}$. In this work, we restrict our attention to the $z$-component of the DMI, leading to
$\color[rgb]{0,0,0}H_{DMz}=D_{z}\sum_{i<j}\left(\sigma_{i}^{x}\sigma_{j}^{y}-\sigma_{i}^{y}\sigma_{j}^{x}\right),$} (8)
where $D_{z}$ quantifies the interaction strength along the $z$-axis.
Experimental and first-principles studies indicate that DMI strengths can vary significantly across materials. For instance, in two-dimensional ternary compounds of the form $MCuX_{2}$ (with $M$ a $3d$ transition metal and $X$ a group-VIA element), the DMI can range from 0 to 15 meV/atom, with $VCuTe_{2}$ exhibiting a value of 15.2 meV/atom [Ga_DMI]. Likewise, in 2D Janus structures such as Fe/Ir(111) thin films, a DMI strength of approximately 1.7 meV has been reported, contributing to skyrmion formation [Dupe_Fe_IR].
Within the quantum-battery setting, the Heisenberg interaction governs the intrinsic spin–spin coupling responsible for cooperative charging, while the inclusion of DMI introduces anisotropy and spin canting, thereby modifying the population dynamics, enhancing quantum correlations [Gurkan2010943, Jafari20111057], and altering the available pathways for energy transfer. These combined effects influence both ergotropy and charging power [Rahman2024], making Heisenberg exchange and DMI crucial components in understanding and optimizing the performance of spin-chain-based quantum batteries.
In the charging protocol, the quantum system is initialized in the ground state of the battery Hamiltonian $H_{B}$. At time $t=0$, the charging process is initiated by abruptly switching to the charging Hamiltonian $H_{C}$. After a time interval $\tau$, the Hamiltonian is switched back to $H_{B}$, thereby completing the charging cycle. To investigate the role of the battery Hamiltonian during the charging dynamics, we introduce a weighted contribution of $H_{B}$ to the charging Hamiltonian [shukla2025]. The time-dependent Hamiltonian is thus defined as
$H(t)=\begin{cases}H_{B},&t<0,\\ H_{C}+\lambda H_{B},&0\leq t\leq\tau,\\ H_{B},&t>\tau,\end{cases}$ (9)
where $\lambda\in[0,1]$ controls the contribution of the battery Hamiltonian during the charging interval. For $\lambda=0$, the dynamics is governed purely by the charging Hamiltonian $H_{C}$, while for $\lambda>0$ the battery Hamiltonian partially contributes during charging.
In this sense, $\lambda$ acts as a continuous switch that interpolates between purely driven dynamics ($\lambda=0$) and dynamics where the battery Hamiltonian partially contributes during charging ($\lambda>0$). This goes beyond a simple ON/OFF protocol and allows us to probe intermediate regimes that are commonly relevant in realistic experimental settings.
Since the Hamiltonian is static during the charging window $0\leq t\leq\tau$, the unitary time evolution operator governing the charging dynamics is given by
$U=e^{-i(H_{C}+\lambda H_{B})t}.$ (10)
Throughout the charging process, the system is assumed to be thermally isolated, and all energy stored in the battery is quantified with respect to the battery Hamiltonian $H_{B}$, which defines the reference energy spectrum used in the computation of ergotropy.
@@
Figure 4: Ergotropy for Ising model ($\delta=1,\Delta=0$) with DM interaction strength D = 0, 1, 2
(a) Power for Ising model ($\delta=1,\Delta=0$) with DM interaction strength D = 0, 1, 2
II.1 Spin chain configurations